### v1

@@ 4.1 A preconditioned correction equation
(4.18) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\sigma_{m+1}-\tau|}\right)^{\left[\frac{j}{2}\right]}$
with $\sigma_{m+1}$ the $(m+1)$th closest singular value to $\tau$ in terms of the labeling rule (1.2).
Proof.
Under the hypothesis, we take $\widetilde{P}_{m}=P_{m}$ and $\widetilde{Q}_{m}=Q_{m}$ in (4.11) with
(4.19) $P_{m}=[u_{m+1},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{m}=[v_{m+1},\dots,v_{N}].$
In this case, the matrix $\widetilde{B}_{m}^{\prime}$ defined in (4.12) is simplified as
(4.20) $B_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&P_{m}^{T}AQ_{m}\\ Q_{m}^{T}A^{T}P_{m}&-\tau I_{N-m}\end{bmatrix}=\begin{bmatrix}-\tau I_{M-m}&\Sigma_{m}^{\prime}\\ (\Sigma_{m}^{\prime})^{T}&-\tau I_{N-m}\end{bmatrix},$
@@
[…]
@@
(4.22) ±$\pm\sigma_{m+1}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
(i) If $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$, then from (4.1) the $m$ singular values $\sigma_{1}$, $\dots$, $\sigma_{m}$ closest to $\tau$ are the $m$ largest ones $\sigma_{\max}$, $\sigma_{\max,2}$, $\dots$, $\sigma_{\max,m}$ of $A$ and the remaining singular values $\sigma_{m+1},\ldots,\sigma_{N}$ are smaller than $\tau$, among which $\sigma_{m+1}=\sigma_{\max,m+1}$. Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (4.22) are negative and lie in
@@
Applying (3.11) to MINRES for (4.10) leads to (4.16).
(ii) If $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$, then the $m$ desired singular values are the $m$ smallest ones $\sigma_{\min}$, $\sigma_{\min,2}$, $\dots$, $\sigma_{\min,m}$ of $A$. In this case, $\sigma_{m+1}=\sigma_{\min,m+1}$. The eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (4.22) lie in
@@
where the negative and positive intervals have the equal length $\sigma_{\max}-\sigma_{\min,m+1}$. Using a similar proof approach to that of (3.14), we obtain (4.17).
(iii) For $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, the $\lambda_{i}$ in (4.22) lie in
(4.23) $[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]\cup[|\tau-\sigma_{m+1}|,\sigma_{\max}-\tau]$
$\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]\cup[|\tau-\sigma_{m+1}|,\sigma_{\max}+\tau]$
Notice that the two intervals in (4.23) have the same length $\sigma_{\max}+\tau-|\tau-\sigma_{m+1}|$. Applying (3.12) to MINRES for (4.10) with this interval union yields (4.18). ∎
Remark 4.
Note that items (i)–(iii) of Theorem 4.2 correspond to $\sigma_{1}=\sigma_{\max},\sigma_{\min}$ and $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$. The convergence rates of MINRES for (4.10) thus depend on
(4.24) $\gamma_{4}=\frac{\tau-\sigma_{\max,m+1}}{\tau+\sigma_{\max,m+1}},\qquad\gamma_{5}=\sqrt{\frac{\sigma_{\min,m+1}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}}},\qquad\gamma_{6}=\frac{|\sigma_{m+1}-\tau|}{\sigma_{\max}+\tau},$
which are considerably bigger than $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (3.21) whenever $\tau-\sigma_{\max,m+1}\gg\tau-\sigma_{\max,2}$, $\sigma_{\min,m+1}-\tau\gg\sigma_{\min,2}-\tau$ and $|\sigma_{m+1}-\tau|\gg|\sigma_{2}-\tau|$, that is, the $m$ clustered singular values $\sigma_{1},\ldots,\sigma_{m}$ are well separated from the others. As a result, MINRES for (4.10) will converge much faster than it does for (2.7) correspondingly.
In practical computations, $\mathcal{R}(\widetilde{U}_{m})\neq\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})\neq\mathcal{R}(V_{m})$. Denote by
(4.25) $\Phi=\angle\left(\mathcal{R}(\widetilde{U}_{m}),\mathcal{R}(U_{m})\right)\quad\mbox{and}\quad\Psi=\angle\left(\mathcal{R}(\widetilde{V}_{m}),\mathcal{R}(V_{m})\right)$
the principal or canonical angle matrices between $\mathcal{R}(\widetilde{U}_{m})$ and $\mathcal{R}(U_{m})$ and between $\mathcal{R}(\widetilde{V}_{m})$ and $\mathcal{R}(V_{m})$, respectively (cf. [4, p. 329]). For $\widetilde{B}_{m}^{\prime}$ in (4.12) and $B_{m}^{\prime}$ in (4.20), we have the following extension of Theorem 3.4.
@@
(4.26) $\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|\leq\|A\|\left(\|\sin\Phi\|+\|\sin\Psi\|\right)^{2},$
where $Z$ is some orthogonal matrix in $\mathbb{R}^{(M+N-2m)\times(M+N-2m)}$; see (4.33).
Proof.
By (4.19), let us decompose the orthonormal $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$, defined in the beginning of section 4.2, as the orthogonal sums
(4.27) $\widetilde{P}_{m}=P_{m}E_{1}+U_{m}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{m}=Q_{m}F_{1}+V_{m}F_{2}$
@@
(4.28) $\|\sin\Phi\|=\|\widetilde{P}_{m}^{T}U_{m}\|=\|E_{2}\|\qquad\mbox{and}\qquad\|\sin\Psi\|=\|\widetilde{Q}_{m}^{T}V_{m}\|=\|F_{2}\|.$
@@
(4.29) $E_{1}=X_{1}\Sigma_{E}X_{2}^{T}\qquad\mbox{and}\qquad F_{1}=Y_{1}\Sigma_{F}Y_{2}^{T}.$
Using the same approach to proving (3.26), we obtain
(4.30) $\|I-\Sigma_{E}\|\leq\|\sin\Phi\|^{2}\qquad\mbox{and}\qquad\|I-\Sigma_{F}\|\leq\|\sin\Psi\|^{2}.$
By the definitions of $P_{m},Q_{m}$ and $U_{m},V_{m}$, we have $U_{m}^{T}AV_{m}=\Sigma_{m}$, $P_{m}^{T}AQ_{m}=\Sigma_{m}^{\prime}$ and $P_{m}^{T}AV_{m}=\bm{0}$, $U_{m}^{T}AQ_{m}=\bm{0}$, where $\Sigma_{m}$ and $\Sigma_{m}^{\prime}$ are defined in (4.2) and (4.21). Therefore, exploiting (4.27) and (4.29), we have
(4.31) $\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}=E_{1}^{T}\Sigma_{m}^{\prime}F_{1}+E_{2}\Sigma_{m}F_{2}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+E_{2}\Sigma_{m}F_{2}.$
Note that $X_{i},Y_{i},i=1,2$ are orthogonal and $\|\Sigma_{m}^{\prime}\|\leq\|A\|$, $\|\Sigma_{m}\|\leq\|A\|$. Exploiting (4.28), (4.30) and (4.31), we obtain
$\|E_{2}\Sigma_{m}F_{2}\|\leq\|A\|\|\sin\Phi\|\|\sin\Psi\|$
and
(4.32) $\|\!\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\!-\!X_{\!2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{\!1}Y_{\!2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\!\Sigma_{m}^{\prime}Y_{\!1}+X_{2}^{T}E_{2}\Sigma_{m}F_{2}Y_{2}\|$
@@
(4.33) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2m)\times(M+N-2m)}.$
By the definitions of $\widetilde{B}_{m}^{\prime}$ in (4.12) and $B_{m}^{\prime}$ in (4.20), we obtain
@@
Then (4.26) follows from applying (4.32) to the above. ∎
Making use of Theorems 4.24.3, we can establish the following convergence results on MINRES for solving (4.10). The proof is analogous to that of Theorem 3.5.
@@
(4.35) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1}+\delta)/(\tau-\sigma_{\max,m+1}-\delta)}}\right)^{j};$
(ii) If $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ and writing $j_{o}=\min\{1,M-N\}$, then
(4.36) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,m+1}-\delta)^{2}-\tau^{2}}}}\right)^{\left[\frac{j-j_{o}}{2}\right]}.$
@@
(4.37) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{m+1}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
The factors in (4.35)–(4.37) are approximations to $\gamma_{4},\gamma_{5}$ and $\gamma_{6}$ in (4.24) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,m+1}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,m+1}-\tau})$ and $\mathcal{O}(\frac{\delta}{|\sigma_{m+1}-\tau|})$, respectively. This indicates that the convergence behavior of MINRES for (4.10) resembles that in the ideal case as $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ is converging. If $\sigma_{1}$ is close to $\sigma_{2}$ but well separated from $\sigma_{m+1}$, MINRES for the preconditioned correction equation (4.10) can converge much faster than it does for the standard correction equation (2.7).
@@ 4.3 Selection strategy of approximate singular triplets in inner preconditioning
The previous results and analysis have shown that it is much better to use MINRES to solve the preconditioned (4.10) rather than (2.7) when there are totally $m$ singular values of $A$ clustered at $\tau$ and their approximate singular vectors already have some accuracy. However, the number $m$ is unknown in practice. We need to set up necessary criteria so as to propose a practical IPJDSVD algorithm.
@@
(4.38) $\frac{|\sigma_{i}-\tau|}{\max\{\sigma_{i},1\}}\leq\textit{pretol}_{1}$
for a sufficiently small $\textit{pretol}_{1}$. Theorems 4.34.4 indicate that if the singular vectors with the $m$ clustered singular values are computed with some accuracy, say $\|\sin\Phi\|+\|\sin\Psi\|\leq\textit{pretol}_{2}$ for an appropriately small $\textit{pretol}_{2}$, then we benefit very much from MINRES for solving the preconditioned (4.10).
During each outer iteration, for the currently computed approximate singular triplet $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,k$, if $\theta_{i}$ and the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ in (2.4) satisfy
(4.39) $|\theta_{i}-\tau|\leq\max\{\theta_{i},1\}\cdot\textit{pretol}_{1}\qquad\mbox{and}\qquad\|r^{(i)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{pretol}_{2},$
then we accept $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ as a reasonably good approximate singular triplet of $A$ and claim $\theta_{i}$ clustered at $\tau$, where $\|A\|_{\mathrm{e}}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$ is a rough estimate of $\|A\|=\sigma_{\max}$ with $\|A\|_{1}$ and $\|A\|_{\infty}$ the $1$- and $\infty$-norms of $A$, respectively. If $A$ is accessible, we calculate $\|A\|_{\mathrm{e}}$ directly; if $A$ is implicit and only products of vectors with $A$ and $A^{T}$ can be computed, we compute $\|A\|_{1}=\max\limits_{1\leq i\leq N}\{\|Ae_{i,N}\|_{1}\}$ and $\|A\|_{\infty}=\max\limits_{1\leq i\leq M}\{\|A^{T}e_{i,M}\|_{1}\}$ with $e_{i,j}$ and the $i$th column of the identity matrix $I_{j}$, $j=N,M$. We take the $\tilde{u}_{i},\tilde{v}_{i}$ to form
(4.40) $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]\qquad\mbox{and}\qquad\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{v}_{i_{\tilde{m}}}],$
where the indices $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ are such that (4.39) is fulfilled, and replace $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$ by them in (4.10).
For $k$ small, there may be no sufficiently good approximate singular triplet that satisfies (4.39). In this case, $\tilde{m}=1$ in (4.40), and (4.10) degenerates to (2.7). However, for $m>1$, as the dimension $k$ increases, more and more computed approximate singular triplets start to converge and fulfill (4.39), so that $\tilde{m}>1$ pairs of approximate singular vectors participate (4.10). Then the inner preconditioning starts to show its strength. In computations, $\tilde{m}$ in (4.40) is determined dynamically according to (4.39). $\tilde{m}$ increases from 1 to $m$, and we have $\tilde{m}=m$ ultimately for $k$ sufficiently large.

### v2

@@ 4.2 Convergence analysis on MINRES for solving (56)
(64) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\sigma_{m+1}-\tau|}\right)^{\left[\frac{j}{2}\right]}$
with $\sigma_{m+1}$ the $(m+1)$th closest singular value to $\tau$ in terms of the labeling rule (2).
Proof 4.3.
Under the hypothesis, we take $\widetilde{P}_{m}=P_{m}$ and $\widetilde{Q}_{m}=Q_{m}$ in (57) with
(65) $P_{m}=[u_{m+1},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{m}=[v_{m+1},\dots,v_{N}].$
In this case, the matrix $\widetilde{B}_{m}^{\prime}$ defined in (58) reduces to
(66) $B_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&P_{m}^{T}AQ_{m}\\ Q_{m}^{T}A^{T}P_{m}&-\tau I_{N-m}\end{bmatrix}=\begin{bmatrix}-\tau I_{M-m}&\Sigma_{m}^{\prime}\\ (\Sigma_{m}^{\prime})^{T}&-\tau I_{N-m}\end{bmatrix},$
@@
(67) $\Sigma_{m}^{\prime}=P_{m}^{T}AQ_{m}=\operatorname{diag}\{\sigma_{m+1},\dots,\sigma_{N}\}\in\mathbb{R}^{(M-m)\times(N-m)}.$
@@
(68) ±$\pm\sigma_{m+1}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
(i) In this case, from (47) the $m$ singular values $\sigma_{1}$, $\dots$, $\sigma_{m}$ closest to $\tau$ are the $m$ largest ones $\sigma_{\max}$, $\sigma_{\max,2}$, $\dots$, $\sigma_{\max,m}$ of $A$, and the remaining singular values $\sigma_{m+1},\ldots,\sigma_{N}$ are smaller than $\tau$, among which $\sigma_{m+1}=\sigma_{\max,m+1}$. Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (68) are negative and lie in
@@
Applying (20) to MINRES for (56) leads to (62).
(ii) In this case, the $m$ desired singular values are the $m$ smallest ones $\sigma_{\min}$, $\sigma_{\min,2}$, $\dots$, $\sigma_{\min,m}$ of $A$, and $\sigma_{m+1}=\sigma_{\min,m+1}$. The eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (68) lie in
@@
where the negative and positive intervals have the equal length $\sigma_{\max}-\sigma_{\min,m+1}$. Using a similar proof approach to that of (23), we obtain (63).
(iii) In this case, the $\lambda_{i}$ in (68) lie in
$[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]$$\cup$ $[|\tau-\sigma_{m+1}|,\sigma_{\max}-\tau]$
(69) $\subset\ [-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]$$\cup$ $[|\tau-\sigma_{m+1}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (69) have the same length $\sigma_{\max}+\tau-|\tau-\sigma_{m+1}|$. Applying (21) to MINRES for (56) with this interval union yields (64).
Remark 4.4.
Note that items (i)–(iii) of Theorem 4.2 correspond to $\sigma_{1}=\sigma_{\max},\sigma_{\min}$ and $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$. The convergence rates of MINRES for (56) thus depend on
(70) $\gamma_{4}=\frac{\tau-\sigma_{\max,m+1}}{\tau+\sigma_{\max,m+1}},\qquad\gamma_{5}=\frac{\sigma_{\min,m+1}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}},\qquad\gamma_{6}=\frac{|\sigma_{m+1}-\tau|}{\sigma_{\max}+\tau},$
which are considerably bigger than $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (30) whenever $\tau-\sigma_{\max,m+1}\gg\tau-\sigma_{\max,2}$, $\sigma_{\min,m+1}-\tau\gg\sigma_{\min,2}-\tau$ and $|\sigma_{m+1}-\tau|\gg|\sigma_{2}-\tau|$, that is, the $m$ clustered singular values $\sigma_{1},\ldots,\sigma_{m}$ are well separated from the others. As a result, MINRES for (56) will converge much faster than it does for (9) correspondingly.
In computations, $\mathcal{R}(\widetilde{U}_{m})\neq\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})\neq\mathcal{R}(V_{m})$. Denote by
(71) $\Phi=\angle\left(\mathcal{R}(\widetilde{U}_{m}),\mathcal{R}(U_{m})\right)\quad\mbox{and}\quad\Psi=\angle\left(\mathcal{R}(\widetilde{V}_{m}),\mathcal{R}(V_{m})\right)$
the canonical angle matrices between $\mathcal{R}(\widetilde{U}_{m})$ and $\mathcal{R}(U_{m})$ and between $\mathcal{R}(\widetilde{V}_{m})$ and $\mathcal{R}(V_{m})$, respectively (cf. [3, p. 329]). For $\widetilde{B}_{m}^{\prime}$ in (58) and $B_{m}^{\prime}$ in (66), we have the following extension of Theorem 3.8.
@@
(72) $\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|\leq\|A\|\left(\|\sin\Phi\|+\|\sin\Psi\|\right)^{2},$
where $Z$ is some orthogonal matrix in $\mathbb{R}^{(M+N-2m)\times(M+N-2m)}$; see (79).
Proof 4.6.
By (65), let us decompose the orthonormal $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$, defined in the beginning of Section 4.2, as the orthogonal sums
(73) $\widetilde{P}_{m}=P_{m}E_{1}+U_{m}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{m}=Q_{m}F_{1}+V_{m}F_{2}$
@@
(74) $\|\sin\Phi\|=\|\widetilde{P}_{m}^{T}U_{m}\|=\|E_{2}\|\qquad\mbox{and}\qquad\|\sin\Psi\|=\|\widetilde{Q}_{m}^{T}V_{m}\|=\|F_{2}\|.$
@@
(75) $E_{1}=X_{1}\Sigma_{E}X_{2}^{T}\qquad\mbox{and}\qquad F_{1}=Y_{1}\Sigma_{F}Y_{2}^{T}.$
Using the same approach to proving (35), we obtain
(76) $\|I-\Sigma_{E}\|\leq\|\sin\Phi\|^{2}\qquad\mbox{and}\qquad\|I-\Sigma_{F}\|\leq\|\sin\Psi\|^{2}.$
By the definitions of $P_{m},Q_{m}$ and $U_{m},V_{m}$, we have $U_{m}^{T}AV_{m}=\Sigma_{m}$, $P_{m}^{T}AQ_{m}=\Sigma_{m}^{\prime}$ and $P_{m}^{T}AV_{m}=\bm{0}$, $U_{m}^{T}AQ_{m}=\bm{0}$, where $\Sigma_{m}$ and $\Sigma_{m}^{\prime}$ are defined in (48) and (67). Therefore, exploiting (73) and (75), we have
(77) $\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}=E_{1}^{T}\Sigma_{m}^{\prime}F_{1}+E_{2}\Sigma_{m}F_{2}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+E_{2}\Sigma_{m}F_{2}.$
Note that $X_{i},Y_{i},i=1,2$ are orthogonal and $\|\Sigma_{m}^{\prime}\|\leq\|A\|$, $\|\Sigma_{m}\|\leq\|A\|$. Exploiting (74), (76) and (77), we obtain
(78) $\|\!\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\!-\!X_{\!2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{\!1}Y_{\!2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\!\Sigma_{m}^{\prime}Y_{\!1}+X_{2}^{T}E_{2}\Sigma_{m}F_{2}Y_{2}\|$
@@
(79) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2m)\times(M+N-2m)}.$
By the definitions of $\widetilde{B}_{m}^{\prime}$ in (58) and $B_{m}^{\prime}$ in (66), we obtain
@@
Then (72) follows from applying (78) to the above.
Making use of Theorems 4.2 and 4.5, we can establish the following convergence results on MINRES for solving (56). The proof is analogous to that of Theorem 3.10.
@@
(81) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\Bigg{(}1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1}+\delta)/(\tau-\sigma_{\max,m+1}-\delta)}}\Bigg{)}^{j};$
(ii) If $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ and writing $j_{o}=\min\{1,M-N\}$, then
(82) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,m+1}-\delta)^{2}-\tau^{2}}}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]}.$
@@
(83) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{m+1}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
The factors in (81)–(83) are approximations to $\gamma_{4},\gamma_{5}$ and $\gamma_{6}$ in (70) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,m+1}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,m+1}-\tau})$ and $\mathcal{O}(\frac{\delta}{|\sigma_{m+1}-\tau|})$, respectively. This indicates that the convergence behavior of MINRES for (56) resembles that in the ideal case as $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ converges. If $\sigma_{1}$ is close to $\sigma_{2}$ but well separated from $\sigma_{m+1}$, MINRES for the preconditioned correction equation (56) can converge much faster than it does for the standard correction equation (9).
@@ 4.3 Selection strategy of approximate singular triplets in inner preconditioning
The previous results and analysis have shown that it is much better to use MINRES to solve the preconditioned (56) rather than (9) when there are totally $m$ singular values of $A$ clustered at $\tau$ and their approximate singular vectors already have some accuracy. However, the number $m$ is practically unknown. We need to set up necessary criteria so as to propose a practical IPJDSVD algorithm.
@@
(84) $\frac{|\sigma_{i}-\tau|}{\max\{\sigma_{i},1\}}\leq\textit{pretol}_{1}$
for a reasonably small $\textit{pretol}_{1}$. Theorems 4.5 and 4.7 indicate that if the singular vectors with the $m$ clustered singular values are computed with some accuracy, say $\|\sin\Phi\|+\|\sin\Psi\|\leq\textit{pretol}_{2}$ for a reasonably small $\textit{pretol}_{2}$, then we can benefit very much from MINRES for solving the preconditioned (56).
During each outer iteration, for the currently computed approximate singular triplet $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,k$, if $\theta_{i}$ and the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ in (6) satisfy
(85) $|\theta_{i}-\tau|\leq\max\{\theta_{i},1\}\cdot\textit{pretol}_{1}\qquad\mbox{and}\qquad\|r^{(i)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{pretol}_{2},$
then we claim $\theta_{i}$ clustered at $\tau$ and accept $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ as a reasonably good approximate singular triplet of $A$, where $\|A\|_{\mathrm{e}}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$ is a rough estimate of $\|A\|=\sigma_{\max}$ with $\|A\|_{1}$ and $\|A\|_{\infty}$ the $1$- and $\infty$-norms of $A$, respectively. We then take the $\tilde{u}_{i},\tilde{v}_{i}$ to form
(86) $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]\qquad\mbox{and}\qquad\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{v}_{i_{\tilde{m}}}],$
where the indices $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ are such that (85) is fulfilled, and replace $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$ by them in (56).
For $k$ small, there may be no sufficiently good approximate singular triplet that satisfies (85). In this case, $\tilde{m}=1$ in (86), and (56) degenerates to (9). However, for $m>1$, as the dimension $k$ increases, more and more computed approximate singular triplets start to converge and fulfill (85), so that $\tilde{m}>1$ pairs of approximate singular vectors participate (56). Then the inner preconditioning starts to show its strength. In computations, $\tilde{m}$ in (86) is determined dynamically according to (85). $\tilde{m}$ increases from 1 to $m$, and we have $\tilde{m}=m$ ultimately for $k$ sufficiently large.